OpenStudy (anonymous):

what is the exact circumference of a circle with a diameter of 29 inches

OpenStudy (anonymous):

Formula of the circumference is \[Circumference = Diameter \times \pi \] Therefore your answer is just: \[Circumference = 29 \pi\]